A spotlight on the ground is shining on a wall 24m

 away. If a woman 2m

 tall walks from the spotlight toward the building at a speed of 1.2m/s,

 how fast is the length of her shadow on the building decreasing when she is 2m

 from the building?


1
Expert's answer
2021-11-16T11:23:27-0500

Solution:


Given "\dfrac{d x}{d t}=1.2 \ m/s"

from graph,

"\begin{aligned}

&\frac{y}{24}=\frac{2}{x} \\

&\Rightarrow y=\frac{48}{x}

\end{aligned}"

differentative w.r.to "\mathrm{t}" ,

"\begin{aligned}

&\Rightarrow \frac{d y}{d t}=48 \frac{d}{d t}\left(\frac{1}{x}\right) \\

&\Rightarrow \frac{d y}{d t}=48\left[-\frac{1}{x^{2}}\right] \frac{d x}{d t} \\

&\Rightarrow \frac{d y}{d t}=48\left[-\frac{1}{(24-2)^{2}}\right](1.2) \ [Using \ (i)]
\\& \Rightarrow \frac{d y}{d t}=-0.119
\end{aligned}"

Thus, the shadow on building is decreasing at a rate of 0.119 m/s.